OpenStudy (anonymous):

Solve the logarithmic equation. Express your solutions in exact form only.. Show all work).. log3(x+10)+log3(x-3)-2=log3(x-2) Solve the logarithmic equation. Express your solutions in exact form only.. Show all work).. log3(x+10)+log3(x-3)-2=log3(x-2) @Mathematics

OpenStudy (anonymous):

u use the formula log(ab)=log a+log b.. log3{(x+10)(x-3)/(x-2)}=2 nd now solve it..

OpenStudy (anonymous):

so (x+10)(x-3)/(x-2)=2^3

jhonyy9 (jhonyy9):

will be x2+7x-30=8x-16 x2 -x -14=0 x_1,_2=(1+/- sqrt(1+56))/2 =(1+/- sqrt57)/2